OpenStudy (anonymous):

i need help finding the limit of the sequence (1 + 1/n) ^ n using l'hopitals rule

OpenStudy (anonymous):

the limit as n goes to infinity I assume?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

I understand taking the ln of it so you get n ln (1+1/n) then you have ln(1+1/n) / 1/n but then i am lost

OpenStudy (anonymous):

you almost did 75% of the problem

OpenStudy (anonymous):

now you're ready to apply the l'hopital's rule, that's taking the derivative of both The numerator and denominator

OpenStudy (anonymous):

so the denominator would be -1/x^2

OpenStudy (anonymous):

you will get: \[\lim_{n \rightarrow \infty} {\ln (1+1/n) \over 1/n}=\lim_{n \rightarrow \infty} { {-1/n^2 \over 1+1/n} \over -1/n^2}=\lim_{n \rightarrow \infty}{1 \over 1+1/n}\]

OpenStudy (anonymous):

which is equal to 1 as you can see.

OpenStudy (anonymous):

so the top is 1 / 1 + 1/n and then you take the derivate of the inside of ln (1 + 1/n) right

OpenStudy (anonymous):

but this is the limit of the function after taking the ln, so the limit of your original function is e^1

OpenStudy (anonymous):

yes you're right!!

OpenStudy (anonymous):

awesome thanks bud

OpenStudy (anonymous):

no problem.. just be a fan ;)

OpenStudy (anonymous):

AnwarA you still there I got another tought one its (1 + 1/sqrt(x))^x

OpenStudy (anonymous):

x is approaching what?

OpenStudy (anonymous):

infinty again

OpenStudy (anonymous):

use the same procedure we did with the first one.

OpenStudy (anonymous):

ok so I got to that part I got ( -1/2x ^ 3/2)/ 1 + 1/sqrt(x) / -1/x^2 but it doesnt look good to simplify to me

OpenStudy (anonymous):

you should use parentheses.